OpenStudy (anonymous):

A freight train car with a mass of 1.2 × 104 kilograms, travelling with a velocity of 23 meters/second, collides with a stationary freight train car of equal mass. What is the momentum of the system after the collision? either: A. 1.2 × 105 kilogram meters/second B. 1.4 × 105 kilogram meters/second C. 2.4 × 104 kilogram meters/second D. 2.8 × 105 kilogram meters/second

OpenStudy (anonymous):

D

OpenStudy (anonymous):



OpenStudy (anonymous):

\[(1200kg)(23 m/s)+(12000kg)(0 m/s) = p_3\] p3 is essentially going to be them both together however the one object had no momentum so all the momentum from the first train has to equal what comes out so. \[12000kg(23 m/s)=27600 kgm/s =27600 N*s=2.8*10^5 N*s\]